OpenStudy (anonymous):

Match each equation in the left column with its solution in the right column. Term: 1. 8(x + 5) = 48 2. 5(2x – 3) = 5 3. 4(x – 8) = 4 4. 3(3x + 1) = 30 Definition: A) x = 1 B) x = 3 C) x = 2 D) x = 9 ~Please help and explain if you can. I just can't seem to wrap my head around how to do it.~

OpenStudy (anonymous):

Just solve each equation and then match

OpenStudy (anonymous):

That's my problem. I can never get the equation right. I always mess up somewhere in the process.

OpenStudy (anonymous):

Or substitute each answer instead of x for example: 8(1+5)=48 8+40=48 So x=1 and so on

OpenStudy (anonymous):

Okay.. So how does (1+5) become 40? That's the part I don't understand.

OpenStudy (anonymous):

Since you have bracket like this: 8(1+5) You multiply the 8 with the numbers in the bracket 8*1=8 and 8*5=40 So since you have the numbers 8 and 40 which are both positive you will get+8+40 which is equal to +48

OpenStudy (anonymous):

Consider this rule so it can help you: \[a(b+c)=ab+ac\]

OpenStudy (anonymous):

You just the multiply the outside number with the numbers in the bracket Till here did you understand?

OpenStudy (anonymous):

Oh. Okay that makes since. Can you help me with second one? Just so I can make sure I've got it?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

We will try solving x

OpenStudy (anonymous):

The equation is: 5(2x – 3) = 5

OpenStudy (anonymous):

Ok so try to get rid of the brackets as I showed you before.

OpenStudy (anonymous):

Ok. Don't I have to move the X first though?

OpenStudy (anonymous):

No Just multiply the outside number with the numbers and variables in the bracket Consider also this rule:\[a(bx+c)=abx+ac\]

OpenStudy (anonymous):

No way that I try to do It is right :/

OpenStudy (anonymous):

OK, Calm down I will help you

OpenStudy (anonymous):

OK 5*2x=?

OpenStudy (anonymous):

And 5 multiply with -3 is equal to -15

OpenStudy (anonymous):

So you have 10x-15=5 Now try solving it

OpenStudy (anonymous):

OH! So you multiply 5 by 2, x, and negative 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

And just to make sure. 4(x – 8) = 4 is D and 3(3x + 1) = 30 is B? Correct?

OpenStudy (anonymous):

If you have solved any equation, you should replace your answer with the variable and perform the calculations to get it. For example replace the 2 you have got with x and try calculating it and if you get 5 which is the answer:5(2(2)-3)=5 then that means that you have solved the equation correctly. Your welcome